OpenStudy (selfconfidence777):

t=c+cb t=total cost rose paid for hat c=price hat b=sales tax (percent) if the hat cost 975$ and the sales tax was 6% , how much did rose paid for the hat?

OpenStudy (anonymous):

t=975+975*6=975+5850+6825

OpenStudy (selfconfidence777):

so i add?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

hat was \(\$795\) and tax is \(6\%\) total price is \[795\times (1.06)\]

OpenStudy (anonymous):

oops \[975\times (1.06)\]

OpenStudy (anonymous):

must be a hell of a hat